I am helping my son with his homework
A=36 mi2,b=__________________,h=12mi


Answer by nyc_function(2741)

What exactly are you looking for?
I assume this is a question about area of a triangle.
The formula is A = 1/2 times base times height or A = (1/2)*b*h
We need to find b or the base.
36 = (1/2)* b * 12
36 = 6b
36/6 = b
6 = b
I hope this helps. If not, tell me specifically what it is your son needs help with.
nycfunction@yahoo.com
Guido